# Estimating the linear relation between variables that are never jointly observed: an application in in vivo experiments

### v1

@@ 6 Discussion
This work focuses on a statistical framework designed for extracting dependencies form in vivo experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed. In this work we chose the basic linear multivariate setting, prioritizing simplicity and computational feasibility. Particularly, the estimator based on the method of moments makes no hypotheses on data distribution and can be calculated explicitly. The estimator based on optimal transport includes a simple optimization problem, and is based on the Gaussian form of the Wasserstein distance but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in whatever case.
However, these approaches are inapplicable in the cases where linear relationship hypothesis cannot be satisfied. For instance, it is the case with predicting survival data with some continuous biomarker, which is of particular interest in research into radio-induced adverse effects. To be able to consider such scenarios, our model can be extended to a more general case, namely with generalized linear model. The optimal transport estimator appears promising in this context given the fact that Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
Another further research direction lies in investigating alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases but require putting priors on distributions. Concerning a likelihood-based approach, the joint likelihood to be maximized can be written in a factorized form due to the independence of $X$ and $Y$: $L(x,y;\beta_{0},\beta_{1},\mu_{x}^{k},\sigma_{\epsilon},\sigma_{x})=L_{X}(x;\mu_{x},\sigma_{x})L_{Y}(y;\beta_{0},\beta_{1},\mu_{x},\sigma_{\epsilon},\sigma_{x})$ for a group $k$ in the univariate setting. The marginal likelihoods can be expressed explicitly, assuming that the model (1) is true, and that $X|(G=k)\sim\mathcal{N}(\mu_{x}^{k},\sigma_{x,k}^{2})$, and thus $Y|(G=k)\sim\mathcal{N}(\beta_{0}+\beta_{1}\mu_{x}^{k},\sigma_{\epsilon}^{2}+\beta_{1}^{2}\sigma_{x,k}^{2})$, with the noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ independent of $X$.
Finally, it would be of interest to work on improving theoretical properties of our estimators, namely correcting the negative bias that appears for both estimators. The latter is particularly important in the case of the optimal transport estimator, which may be arising naturally with the Wasserstein distance (Wasserman,, 2023). Correcting this bias would considerably improve the estimator, making it competitive with the approaches mentioned earlier that make numerous assumptions on the data.
Funding
This work is supported by the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and by Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS.
@@
Acknowledgements: The authors warmly thank Dr. Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
@@ Appendix B Proofs
First note that the assumptions ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$ ensure the existence of $\sigma_{Y}^{2}$ and $\boldsymbol{\Gamma}_{X}$. From the law of large numbers, we have that for all $k\in\{1,\ldots,K\},\ \widehat{\boldsymbol{\mu}}_{1,X}^{k}\to\boldsymbol{\mu}_{X}^{k}$ and $\widehat{\mu}_{Y}^{k}\to\mu_{Y}^{k}$ in probability when $n_{\min}$ tends to infinity. We deduce from the continuous mapping theorem that $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ and $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}$ in probability. Under hypothesis $\mathbf{H}_{1}$, the inverse being continuous in a neighborhood of $\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ another application of the continuous mapping theorem gives that $\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\to\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}$ and
@@
in probability as $n_{\min}$ tends to infinity. The law of large numbers gives that $\widehat{\boldsymbol{\Gamma}}_{X}^{2}\to\boldsymbol{\Gamma}_{X}^{2}$ and $\widehat{\sigma}_{Y}^{2}\to\sigma_{Y}^{2}$ in probability and we deduce, with another application of the continuous mapping theorem, that $\widehat{\sigma}_{Y}^{2}-\widehat{\boldsymbol{\beta}}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{2}\widehat{\boldsymbol{\beta}}\to\sigma_{Y}^{2}-\boldsymbol{\beta}^{\top}\boldsymbol{\Gamma}_{X}^{2}\boldsymbol{\beta}=\sigma_{\epsilon}^{2}$ in probability as $n_{\min}$ tends to infinity. ∎
@@
The proof is based on Lemma 2.9 and Theorem 2.1 in Newey and McFadden, (1994), which are recalled in Appendix A. The law of large numbers and the continuous mapping theorem give us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity. Consider now $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$. We have,
@@
with Cauchy-Schwarz inequality and $A_{n,k}=\mathcal{O}_{p}(1)$ because $\|\widehat{\boldsymbol{\mu}}_{X}^{k}\|=O_{p}(1)$, $\widehat{\mu}_{Y}^{k}=O_{p}(1)$ and for some constant $C_{1}$ that does not depend on $\boldsymbol{\alpha}$ and $\boldsymbol{\gamma}$, $\|\boldsymbol{\alpha}+\boldsymbol{\gamma}\|\leq C_{1}<\infty$ because $\Theta$ is supposed to be compact. On the other hand, we have
@@
$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (8)
@@
$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (9)
where $B_{n}^{k}=O_{p}(1)$. Combining previous inequalities, we get
$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (10)
@@
$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (17)
@@
$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (21)
Since model (1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (17) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution. As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$ and by the continuous mapping theorem, $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability. It can also be noted that, under the moment condition ${\mathbb{E}}\left[\|\mathbf{X}\|^{4}|G=k\right]<\infty$, the central limit theorem gives that $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}-\boldsymbol{\Gamma}_{X}^{k}\right)$ converges in distribution to a centered Gaussian multivariate distribution, and we deduce with the Cramer-Wold device, the continuous mapping theorem and Slutsky’s theorem that the second component of $\nabla\varphi_{n}$ multiplied by $\sqrt{n}$ also in distribution to a centered Gaussian random vector. It is immediate to deduce that the same convergence result holds for the third component, which is to say that $\sqrt{n}\left(\sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\right)$ converges in distribution to a centered Gaussian random variable. We finally deduce, with the Cramer-Wold device, that (iii) is fulfilled. To prove that (iv) also holds, consider the Hessian matrix of functional $\varphi_{n}$, evaluated at $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$:
@@
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (22)
@@
$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (28)
where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (28) is strictly positive. We now have to show that
$\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}=2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}-2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}$ (31)
$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (32)
@@
using the fact that $\sum_{k}(\sqrt{\pi_{k}})^{2}=1$. It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which can not happen under hypothesis $\mathbf{H}_{1}$. The second part at the righthand side of (32) is handled the same way. We have
@@
and consequently the determinant of $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ is strictly positive. To finish the proof, it remains to check that in a neighborhood $\mathcal{N}$ of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, we have

### v3

@@ 9 Discussion
This work focuses on a statistical framework designed for extracting dependencies from experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed. In this work we chose the basic linear multivariate setting, prioritizing simplicity and computational feasibility. Particularly, the estimator based on the method of moments makes no hypotheses on data distribution and can be calculated explicitly. The estimator based on optimal transport includes a simple optimization problem, and is based on the Gaussian form of the Wasserstein distance but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in whatever case.
However, these approaches are inapplicable in the cases where linear relationship hypothesis cannot be satisfied. For instance, it is the case with predicting survival data with some continuous biomarker, which is of particular interest in research into radio-induced adverse effects. To be able to consider such scenarios, our model can be extended to a more general case, namely with generalized linear model. The optimal transport estimator appears promising in this context given the fact that the Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
Another further research direction lies in investigating alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases but require putting priors on distributions.
Finally, it would be of interest to work on improving theoretical properties of our estimators for finite samples, namely correcting the negative bias that appears for both estimators. The latter is particularly important in the case of the optimal transport estimator, which may be arising naturally with the Wasserstein distance (Wasserman,, 2023). Correcting this bias would considerably improve the estimator, making it competitive with the approaches mentioned earlier that make numerous assumptions on the data.
Software availability
Python code and the experimental data are freely available at github.com/parsenteva/vivo_lm.
Acknowledgements
The authors gratefully acknowledge the funding from the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and from the Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS. The authors also warmly thank Dr. Agnès François and Dr. Annaïg Bertho for the experimental data, as well as Dr. Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
@@ Appendix B Proofs
First note that the assumptions ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$ ensure the existence of $\sigma_{Y}^{2}$ and $\boldsymbol{\Gamma}_{X}$. From the law of large numbers, we have that for all $k\in\{1,\ldots,K\},\ \widehat{\boldsymbol{\mu}}_{1,X}^{k}\to\boldsymbol{\mu}_{X}^{k}$ and $\widehat{\mu}_{Y}^{k}\to\mu_{Y}^{k}$ in probability when $n_{\min}$ tends to infinity.
We deduce from the continuous mapping theorem that $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ and $\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}\to\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{Y}$ in probability. Under hypothesis $\mathbf{H}_{1}$, the inverse being continuous in a neighborhood of $\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}$ another application of the continuous mapping theorem gives that $\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\to\left(\boldsymbol{\mu}_{1,X}^{\top}\mathbf{w}\boldsymbol{\mu}_{1,X}\right)^{-1}$ and
@@
in probability as $n_{\min}$ tends to infinity.
The law of large numbers gives that $\widehat{\boldsymbol{\Gamma}}_{X}^{2}\to\boldsymbol{\Gamma}_{X}^{2}$ and $\widehat{\sigma}_{Y}^{2}\to\sigma_{Y}^{2}$ in probability and we deduce, with another application of the continuous mapping theorem, that $\widehat{\sigma}_{Y}^{2}-\widehat{\boldsymbol{\beta}}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{2}\widehat{\boldsymbol{\beta}}\to\sigma_{Y}^{2}-\boldsymbol{\beta}^{\top}\boldsymbol{\Gamma}_{X}^{2}\boldsymbol{\beta}=\sigma_{\epsilon}^{2}$ in probability as $n_{\min}$ tends to infinity. ∎
@@
The proof is based on Lemma 2.9 and Theorem 2.1 in Newey and McFadden, (1994), which are recalled in Appendix A.
The law of large numbers and the continuous mapping theorem give us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\Theta$, $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity.
Consider now $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$. We have,
@@
with Cauchy-Schwarz inequality and $A_{n,k}=\mathcal{O}_{p}(1)$ because $\|\widehat{\boldsymbol{\mu}}_{X}^{k}\|=O_{p}(1)$, $\widehat{\mu}_{Y}^{k}=O_{p}(1)$ and for some constant $C_{1}$ that does not depend on $\boldsymbol{\alpha}$ and $\boldsymbol{\gamma}$, $\|\boldsymbol{\alpha}+\boldsymbol{\gamma}\|\leq C_{1}<\infty$ because $\Theta$ is supposed to be compact.
On the other hand, we have
@@
$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (B.1)
@@
$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (B.2)
where $B_{n}^{k}=O_{p}(1)$.
Combining previous inequalities, we get
$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (B.3)
@@
$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (B.4)
@@
$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (B.5)
Since model (2.1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (B.4) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution. As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$ and by the continuous mapping theorem, $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability. It can also be noted that, under the moment condition ${\mathbb{E}}\left[\|\mathbf{X}\|^{4}|G=k\right]<\infty$, the central limit theorem gives that $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}-\boldsymbol{\Gamma}_{X}^{k}\right)$ converges in distribution to a centered Gaussian multivariate distribution, and we deduce with the Cramer-Wold device, the continuous mapping theorem and Slutsky’s theorem that the second component of $\nabla\varphi_{n}$ multiplied by $\sqrt{n}$ also in distribution to a centered Gaussian random vector. It is immediate to deduce that the same convergence result holds for the third component, which is to say that $\sqrt{n}\left(\sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\right)$ converges in distribution to a centered Gaussian random variable. We finally deduce, with the Cramer-Wold device, that (iii) is fulfilled.
To prove that (iv) also holds, consider the Hessian matrix of functional $\varphi_{n}$, evaluated at $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$:
@@
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (B.6)
@@
$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (B.7)
where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (B.7) is strictly positive. We now have to show that
$\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}=2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}-2\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)\left(\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}\right)^{\top}$
$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (B.8)
@@
using the fact that $\sum_{k}(\sqrt{\pi_{k}})^{2}=1$. It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which can not happen under hypothesis $\mathbf{H}_{1}$. The second part at the righthand side of (B.8) is handled the same way. We have
@@
and consequently the determinant of $\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ is strictly positive.
To finish the proof, it remains to check that in a neighborhood $\mathcal{N}$ of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$, we have